Calculate 1) GDP at MP

(Rs. '000)

2,000

40

560

500

100

60

200

60

2) NNP at FC

1. Consumption of fixed capital

2. Employer's contribution to social security schemes

3. Rent

4. Interest

5. Profits

6. Royalty

7. Wages and salaries

8. Net indirect taxes

9. Net factor income from abroad

(Crores)

34

30

10

20

25

5

170

38

(-) 3


1
Expert's answer
2022-01-06T10:04:55-0500

(1) To get the GDP at MP given the data

Using the Income approach:

The GDP at MP is there by calculated by

GDPMP=(employer's contribution to social

security) + (Rent) + (Interest) +

(Profit) + ( Consumption of fixed

capital) - (Net indirect tax)

= 2000 + 40 + 560 + 500 + 1000 - 200

= 3900

Hence, the GDPMP is 3900 Rs


(2) To calculate the NNP at FC given the formula

NNPMp = GDPMP - Consumption of fixed

capital + Net factor income from

abroad - Net direct taxes


Now, we must first find the GDPMP

GDPMP = (employer's contribution to social

security) + (Rent) + (Interest) +

(Profit) + ( Consumption of fixed

capital) - (Net indirect tax)

= 30 + 10 + 20 + 25 + 34 - 38

= 81

"\therefore" NNPMP = 81 - 34 + (-3) - 38

= 81 - 34 - 3 - 38

= 6

Hence, the NNPMP is 6